OpenStudy (anonymous):

Let f(x) = 3x+1 and g(x) = x2 Part 1 [2 points] Find f(g(x)) Part 2 [2 points] Find g(f(x)) Part 3 [4 points] Use complete sentences to explain any difference in part 1 and part 2.

OpenStudy (oaktree):

Part 1:

OpenStudy (oaktree):

f(x^2) = 3x^2+1

OpenStudy (oaktree):

g(3x+1) = (3x+1)^2 = 9x^2+6x+1

OpenStudy (oaktree):

And Part 3:

OpenStudy (oaktree):

You can see the difference yourself!